OpenStudy (anonymous):

HELP FAN/MEDAL FOR FIRST RIGHT ANSWER FAST THANK YOU Giselle is going to frame a portrait of the family and place it on the mantle in the family room. The portrait is 10 inches longer than it is tall and will take up a total area of 1344 square inches once it is inside a frame that adds 2 inches to the portrait's width and height. Find the dimensions and area of the unframed portrait.

OpenStudy (solomonzelman):

The portrait is 10 inches longer than it is `tall ` grammar ?? :P

OpenStudy (solomonzelman):

1344 is including the 2 inch frame or not ?

OpenStudy (anonymous):

I dont know..

OpenStudy (solomonzelman):

I would think that 1344 is including the frame, and you are asked to find the area not including the frame.

OpenStudy (solomonzelman):

1344 = ℓ × w 1344 = ℓ × (ℓ +10) 1344 = ℓ² +10l completing the square is the fastest, sorry for the shortcut that I'll use 25+1319 = ℓ² +10l 1319 = ℓ² +10ℓ+25 1319 = (ℓ + 5)² 36.32= ℓ + 5 not taking the negative answer, because the side of portrait can only be positive. 31.32= ℓ So the dimensions with the frame are 31.32 and 41.32